A physics instructor, in a lab, charged an electrostatic generator to 25 μC, then carried it into the lecture hall. The net electric flux in Nm2/C through the lecture hall walls is:
1
Expert's answer
2020-11-27T14:03:12-0500

Solution

According to gauss law flux through the closed volume is given by

"\phi=\frac{\Sigma Q}{\epsilon_0}"

Here

"\phi=\frac{ 25\times10^{-6} }{8.85 \times10^{-12} }=2.82\times 10^6" N-m2/C